OpenStudy (anonymous):

Two students in your class , Hunter and Maggie are disputing a function. Hunter says that for the function , between x=-2 and x=2 the average rate of change is 0. Maggie says that fir the function between, x=-2 and x=2, the graph goes through a turning point and back down. Explain how Hunter and Maggie can both be correct, using complete sentences.

jimthompson5910 (jim_thompson5910):

Were you able to get anywhere with this?

OpenStudy (anonymous):

Nope

jimthompson5910 (jim_thompson5910):

The average rate of change formula is \[\Large A = \frac{f(x_1)-f(x_2)}{x_1-x_2}\] If A = 0, then \(\Large f(x_1)-f(x_2) = 0\) which means that \(\Large f(x_1) = f(x_2)\)

jimthompson5910 (jim_thompson5910):

So if the average rate of change from x = -2 to x = 2 is 0, then f(-2) = f(2)

jimthompson5910 (jim_thompson5910):

Now if "the graph goes through a turning point and back down", then we know that f(-2) starts off low and as x increases, y increases. This happens until we hit a turning point (let's say in the midpoint) and then come back down to x = 2 So it might look something like this